40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14826     Entry time: Sun Aug 4 14:39:41 2019
Author: gautam 
Type: Update 
Category: General 
Subject: some lab activity 
  1. Unresponsive c1psl, c1iool0, c1auxey and c1iscaux VME crates were keyed.
  2. c1psl channels were burt-restored, did a burtrestore, and re-locked the PMC. Tweaked the pointing into the PMC on the PSL table to increase the PMC transmission from ~0.69 to ~0.71.
  3. Re-locked IMC. Ran WFS offset script to relieve the ~100 DAC counts (~10 urad) DC offset from the WFS servos to the IMC suspensions (a serious calibration of this into physical units should be made part of the planned 40m WFS activity). Now that I think about it, since we change the IMC alignment to match the input beam alignment, some post-IMC clipping could modulate the power incident on the ITMs, which is a source of error for the arm cavity loss determination using DC reflection. We need a better normalizing data stream than the IMC transmission.
  4. The IFO_OVEREVIEW medm screen was modified such that the threshold for the PMC transmitted beam to be visible was lowered from 0.7 to 0.6, so that now there is a continuous beam line from the NPRO to the PRM when the IMC is locked even when the PMC transmission degrades by 5% due to thermally driven pointing drifts on the PSL table.
  5. The wmctrl utility on pianosa wasn't working so well, I wasn't able to use my usual locking MEDM autoconfig scripts. Turned out to be due to a zombie MEDM window which I killed with xkill, now it is working okay again.
  6. The misaligned XARM was re-aligned and the loss measuring PDA520 at the AS port was removed from the beam path (mainly to avoid ADC saturations the fringing Michelson will cause).
  7. I noticed that the ETMX Oplev HeNe SUM level has degraded to ~50% of its power level from 200 days ago [Attachment #1], may need a new HeNe here soon. @Chub, do we have spare HeNes in stock?

I want to collect some data with the arms locked to investigate the possibility/usefullness of having seismic feedforward implemented for the arms (it is already known to help the IMC length and PRC angular stability at low frequencies). To facilitate diagnostics I modified the file /users/Templates/Seismic/Seismic_vs_TRXTRYandMC.xml to have the correct channel names in light of Lydia's channel name changes in 2016. Looking at the coherence data, the alignment of the cartesian coordinate system of the Seismometers at the ends and the global interferometer coordinate system can be improved.

I don't know if for the MISO filter design if there is any difference in using TRX/TRY as the target, or the arm length control signal.

Data collection started at 1249018179. I've setup a script running in a tmux shell to turn off the LSC enable in 2 hours.

Attachment 1: ETMX_OL.png  30 kB  Uploaded Sun Aug 4 16:20:46 2019  | Hide | Hide all
ETMX_OL.png
Attachment 2: Seismic_TRXTRYandMC_4Aug2019.pdf  118 kB  Uploaded Sun Aug 4 16:44:30 2019  | Hide | Hide all
Seismic_TRXTRYandMC_4Aug2019.pdf
ELOG V3.1.3-